News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 28 July 2006 8.00am EDT

During the last week the observing schedule was replanned to include an observation of PKS 2155-304 that was accepted as a one day turn-around Target Of Opportunity (TOO) on July 27. Loads were replanned and uplinked on July 28. An observation of M17 was impacted by the TOO and will be rescheduled in a later load. Observations of SDS J1235+5910, Maffei 2 and A478 were added to the schedule as a result of the replan.

Several real-time procedures were conducted during last week associated with a flight software patch to add on-board monitoring of selected propulsion line and valve temperatures. On 25 July a patch was uplinked that replaced 21 of the Liquid Apogee Engine temperature readings (unused since the ascent phase) with 21 -Z side propulsion line and valve temperatures. A real-time procedure was performed during each contact with Chandra following the patch uplink to verify proper operation of the patch. On 27 July a patch was uplinked that implemented the new on-board monitor. Chandra has experienced decreased thermal margins as the mission has proceeded and the new on-board monitor will mitigate the risk of a frozen propulsion line by ensuring a transition to normal sun mode in the event that temperatures fall below a safe trigger threshold.

A realtime procedure was completed on July 27 to update the on-board Ephemeris.

The schedule of targets for next week is shown below and includes an observation of M31 coordinated with HST.

------------------------------------------
NGC 1507                ACIS-S       Jul 30
M31                     ACIS-I
1RXS J162526.9+455      ACIS-S
XMMU J132218.3-164      ACIS-S
AX J1733.9-3112         ACIS-I       Jul 31
Radiation Belts
M17 Pointing I          ACIS-I
M87                     ACIS-S       Aug 2
Radiation Belts
Kepler's Supernova      ACIS-S       Aug 3
Radiation Belts                      Aug 5
NGC 720                 ACIS-S       Aug 6
M17 Pointing I          ACIS-I

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports